OpenStudy (anonymous):

can someone help show me how this is true... u'(t) = -sin^2(t)/cos(t) = -sect + cost

myininaya (myininaya):

?

myininaya (myininaya):

I don't think I understand what your asking. Are you asking does -(sint)^2/cost simplify to -sect+cost

OpenStudy (anonymous):

yes

myininaya (myininaya):

I honestly don't think it does equal

OpenStudy (anonymous):

thats how i feel, it is in the solution manual and in order to get the same answer I need to convince myself that it is true...

myininaya (myininaya):

what happens if you plug zero in on both sides? the equation doesn't hold

myininaya (myininaya):

oops it does hold for t=0

myininaya (myininaya):

let me think some more

myininaya (myininaya):

omg im moron i got it

OpenStudy (anonymous):

yahh!

myininaya (myininaya):

okay (sint)^2+(cost)^2=1 so (sint)^2=1-(cost)^2 therefore -(sint)^2=(cost)^2-1 => -(sint)^2/cost=[(cost)^2-1]/cost

myininaya (myininaya):

split those fractions up! :)

myininaya (myininaya):

so we have (cost)^2/cost - 1/cost= cost-sect= -sect+cost

myininaya (myininaya):

any questions? :)

OpenStudy (anonymous):

good eye thanks!

myininaya (myininaya):

sometimes im so slow

OpenStudy (anonymous):

haha saved me a headache

OpenStudy (anonymous):

:)